The sum of the square of a positive number and the square of 3 more than the number is 29.can you show steps how to work this out and the correct answer ?
Found 2 solutions by JThomson, ikleyn:
Answer by JThomson(12)

Let the number be 'x'

The square of the number is x^2

3 more than the number is (x+3);

The square of 3 more than the number is (x+3)^2

So x^2+(x+3)^2 = 29

x^2+(x^2+6x+9)-29=0

2x^2+6x-20=0

x^2+3x-10=0

(x+5)(x-2)=0

x=-5 and 2

But since x must be a positive number, x=2 .


Answer by ikleyn(52946)

.
Let the number be 'x'

The square of the number is x^2

3 more than the number is (x+3);

The square of 3 more than the number is (x+3)^2

So x^2+(x+3)^2 = 29

x^2+(x^2+6x+9)-29=0

2x^2+6x-20=0

x^2+3x-10=0

(x+5)(x-2)=0

x=-5 and 2

But since x must be a positive number, x=2.

Answer. The number is 2.